OpenStudy (anonymous):

I need to find the flux through a disk of radius 1 centered at (1,1,1) with its normal pointing towards the origin. How do i parameterize this disk?

OpenStudy (anonymous):

I found that it lies in the plane -x-y-z=3

OpenStudy (anonymous):

dont parametrize it

OpenStudy (anonymous):

should i take the intersection between this plane and a sphere centered at 111

OpenStudy (anonymous):

find a vec tor normal to it, with magnitude of the area of the disk

OpenStudy (anonymous):

-1,-1,-1

OpenStudy (anonymous):

whats the vec tor field??

OpenStudy (anonymous):

f=<z,x,y>

OpenStudy (anonymous):

I cant use divergence theorem right?

OpenStudy (anonymous):

im thinking

OpenStudy (anonymous):

what with 0 volume.

OpenStudy (anonymous):

thats whts confusing me..no i dont think divergence thm is applicabe..old way then

OpenStudy (anonymous):

yeah so i need a parametric surface

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thats where im stuck

OpenStudy (anonymous):

thats wht i really dunno..;.parametrizing surfaces

OpenStudy (anonymous):

i can find the intersection between the sphere and a plane but its an ugly integral

OpenStudy (anonymous):

I know theres another way to do it using tangents to the normal but i cant remember

OpenStudy (anonymous):

wht u do

OpenStudy (anonymous):

well the disk lies in the plane -x-y-z=3 and theres a sphere with rad 1 cestered at (1.1.1) of the form <rcos(theta)sin(phi), rsin(theta)sin(phi), rcos(phi)>

OpenStudy (anonymous):

go on.bt watll u do wid d sphere?

OpenStudy (anonymous):

Im trying to figure it out i cant remember exactly

OpenStudy (anonymous):

i know paramet for cylinders and all bt nt this

OpenStudy (anonymous):

Using vectors, generally if t is the parameter then and point P on the circle is given by; LaTeX Code: P = R\\cos(t) \\vec{u} + R\\sin(t) \\;\\;\\vec{n}\\times\\vec{u} + c

OpenStudy (anonymous):

check this out: http://www.physicsforums.com/showthread.php?t=123168

OpenStudy (anonymous):

you need 2 parameters for surfaces

OpenStudy (anonymous):

go down to the post where it writes a parametric eqn for a circle in 3d..our surface is actually a circle at some angle

OpenStudy (anonymous):

read it up

OpenStudy (anonymous):

uve got t and the vector u here as parameters

OpenStudy (anonymous):

u add u the components of i,j,k

OpenStudy (anonymous):

kool ill try the eq on that link

OpenStudy (anonymous):

u get x,y,z in terms of t and u

OpenStudy (anonymous):

i think thatll parametrize it

OpenStudy (anonymous):

U is an orthognal vector not a parameter

OpenStudy (anonymous):

t and r are the parameters

OpenStudy (anonymous):

r is the radius mate...u is the parametric vector from the centre to that point and n is the orthogonal vector

OpenStudy (anonymous):

can you use polar coords?

OpenStudy (anonymous):

you have to integrate over the radius to get a disk

OpenStudy (anonymous):

thats how far i cn get at parameterizing it

OpenStudy (anonymous):

i g2g.. i learnt a lot thru this discusion..thnx..hope u get it

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

polar is the way to go, I'm not sure what him1618 meant

OpenStudy (anonymous):

its in 3 space so its in cylindrical

OpenStudy (anonymous):

just have z=z be one condition

OpenStudy (anonymous):

if z doesn't vary, well i guess cylindrical still applies, anyway the jacobian is still r

OpenStudy (anonymous):

but z does change the disk is tilted in 3 dimentions